OpenStudy (anonymous):

During a dry month, it rained 2 1/3 inches, which is 7 2/3 inches less than the average. Consider the subtraction equation you would solve for the average monthly rainfall. What operation must be done on both sides to solve the equation? Add 2 1/3 Subtract 2 1/3 Add 7 2/3 subtract 7 2/3

OpenStudy (anonymous):

Let x be the average monthly rainfall. So, you would get the equation: 2 1/3+7 2/3=x What do you have to do to find x?

OpenStudy (anonymous):

x - 7 2/3 = 2 1/3 That's how they want you to set it up

OpenStudy (anonymous):

To solve that ... what do you do to both sides?

OpenStudy (anonymous):

This is an absolutely ridiculous question though, because no reasonable person would use algebra to solve this. Just add 7 2/3 and 2 1/3 to get the answer.

OpenStudy (anonymous):

x - 7 2/3 = 2 1/3 To get rid of the 7 2/3....add 7 2/3 to both sides.

OpenStudy (anonymous):

Ok